# DMAligner: Enhancing Image Alignment via Diffusion Model Based View Synthesis

### v1

@@ 5 Conclusion
This work represents a significant paradigm shift in image alignment by introducing a generation-based solution. The proposed DMAligner framework, a diffusion-based approach for image alignment through alignment-oriented view synthesis, effectively addresses challenges such as ghosting artifacts, occlusions, and illumination changes that previous methods struggle with. Notably, a Dynamics-aware Mask Producing (DMP) module is introduced to enable the diffusion model to better handle dynamic foreground regions that are prone to errors. Additionally, the Dynamic Scene Image Alignment (DSIA) dataset is developed specifically for training. Extensive experiments demonstrate that DMAligner outperforms previous methods with unmatched performance and exhibits strong cross-domain capabilities, excelling in both synthetic and real-world scenarios.

### v2

@@ 5 Conclusion
This work represents a significant paradigm shift in image alignment by introducing a generation-based solution. The proposed DMAligner framework, a diffusion-based approach for image alignment through alignment-oriented view synthesis, effectively addresses challenges such as ghosting artifacts, occlusions, and illumination changes that previous methods struggle with. Notably, a Dynamics-aware Mask Producing (DMP) module is introduced to enable the diffusion model to better handle dynamic foreground regions that are prone to errors. Additionally, the Dynamic Scene Image Alignment (DSIA) dataset is developed specifically for training. Extensive experiments demonstrate that DMAligner outperforms previous methods with strong cross-domain capabilities, excelling in both synthetic and real-world scenarios.
Figure 8: Stress test: performance vs. motion magnitude.
Acknowledgements. This work was supported by Sichuan Science and Technology Program under grant No. 2025ZYD0181, in part by the National Natural Science Foundation of China (NSFC) under grant Nos. 62372091 and 62402402, in part by the Hainan Province Science and Technology Plan Project under Grant ZDYF2024(LALH)001, and in part by the Fundamental Research Funds for the Central Universities under grant No. 2682025CX116.